OpenStudy (matthew120):

What is the distance between the points (5, –6) and (5, 2)? A. 1 B. 3 C. 4 D. 8

OpenStudy (anonymous):

Hi

OpenStudy (anonymous):

Link please?

OpenStudy (matthew120):

sup there no pic

OpenStudy (anonymous):

oh lol

OpenStudy (anonymous):

c is the answer

OpenStudy (anonymous):

one sec

OpenStudy (anonymous):

plz medel

OpenStudy (anonymous):

No thats wrong

OpenStudy (anonymous):

|-6|-2 is 4 so the answer is 4

OpenStudy (anonymous):

u have to use absolute value to solve this. The distince from -6 to 2 is 4. u need to ignore the x axsis because it is the same dont listen to this guy he dosn't know what he is doing

OpenStudy (matthew120):

hey dont fight

OpenStudy (anonymous):

may i have the medel plz?

OpenStudy (anonymous):

Well if you do something to one side of the equasion you have to do it to ther so... I-6I=6 now do it to -2.... I-2I=2 so do da math now

OpenStudy (matthew120):

i just want the answer

OpenStudy (anonymous):

6 + 2 = 8 D is answer

OpenStudy (matthew120):

i wiil look up the question to see whos right

OpenStudy (anonymous):

D

OpenStudy (matthew120):

@BillyBob12345678910

OpenStudy (anonymous):

He left

OpenStudy (anonymous):

yay so i got it right

OpenStudy (matthew120):

billy is right

OpenStudy (anonymous):

dude u got it wrong! in ur face sucker